A motorboat traveled for 1 hour with a 2 mph current. The return trip against the same current took 2 hours. Find the speed of the motorboat in still water.
Answer by TimothyLamb(4379)

x = speed of boat in still water
y = one-way trip distance
c = speed of current = 2 mph
---
s = d/t
---
downstream:
x + c = y/1
x + 2 = y/1
---
upstream:
x - c = y/2
x - 2 = y/2
---
put the system of linear equations into standard form
---
x + 2 = y/1
x - 2 = y/2
---
x - y/1 = -2
x - y/2 = 2
---
x - y = -2
x - 0.5y = 2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = speed of boat in still water = 6 mph
y = one-way trip distance = 8 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php